Eng Half Night Plan UT 27 May 2019

  • TCS 2019A test (regardless of MODS-2 IMCS status)
  • MODS-2 IMCS laser repair did not succeed
Decision Tree:

If MODS-2 IMCS laser is repaired (or just doing MODS-1 checks)
  1. Imaging Skyflats at twilight (if clear)
  2. On-sky collimation check - 0.25 hr
  3. Shivaei Payback Program - AZ - 2hrs
    • or
  4. Rotator Center check - 0.5hr
  5. Rotator ZP check with astrometric masks - 0.5 hr
  6. Transform collection / verification - 1hr
  7. Dither test - 0.5-1hr
  8. Spectrophotometic Star measurment - 0.5 hr (if clear)
  9. Photometric Standard star collection - 0.5-1hr (if clear)
  10. Poor weather E task:
    • Examine behavior of open-loop tracking near zenith with MODS.
if MODS-2 IMCS laser is not ready
  1. Still, it could be useful to do some MODS checkout that does not depend on the flexure compensation:
    1. MODS twilight flats. MODS2 will not be too good so I'll put in the title that they were taken without flexcomp.
    2. MODS collimation
  2. LUCI-1 rotator ZP check - need to use alternate method to mask - 0.5hr
  3. LUCI-1/AGW-1 transform data collection (take bino data - DX for "free") - 2.5 hours (update PCSconfig with new rotator ZP then take transform - replace old ZP afterwards)
  4. Poor weather backup program - any amount of time - targets cover entire night
    • H-band library - K and M stars
    • H-band library - case example of QSO rest-frame 1.6micron CO-bandhead - filler
    • Files located at: /home/observer/Service_Observing/E_Tasks_May2019/LUCI_Tasks/ (remote room)
  5. LUCI-1/LUCI-2 Non-linearity on-sky tests - 0.5hr - Requires photometric conditions (seeing not critical)
    1. RA 17hr (IC 4665)
    2. Files located at: /home/observer/Service_Observing/E_Tasks_May2019/LUCI_Tasks/
  6. Poor weather MODS-E task:
    • Examine behavior of open-loop tracking near zenith with MODS.

<h2 NightLog UT 27 May 2019

OSA: J. Williams ISA: O. Kuhn SW: P. Kubanek SO: S. Walsh

Clear skies. Winds borderline and gusting above limits at times. Starting with binoMODS.

01:45 SieveSnap. MODS1 ok, MODS2 (Blue actually locked, with ~200 c above bias, but Red did not. Red counts ~bias)

01:50 Petr testing 2019A. Getting ready to try a MODS preset. Shutter doors went into alarm upon startup up 2019A. Petr on phone with Dan C.

02:12 Ran up LUCI1 and LUCI2. Initialized all for both LUCIs. After the LUCI2 init, the MOS state came up as unknown. I informed Dave; he is working on it.

02:58 Sending a closed-dome active MODS preset to Stone I1.
acqMODS --mods1 --bino -p StoneI1.acq.txt
acqMODS --mods2 --bino -p StoneI1.acq.txt
Worked but of course failed to find a guide star, a "successful failure".

03:02 Josh sending a preset from his side now.

03:07 Olga forgot to put MODS into observing mode. She's done that now, so let's try again for a "successful failure". OK, got it.

03:02 DIMM is reporting 0.0 and not -1.0. Petr says this is a reflective memory issue and that restarting DIMM will fix this.

04:35 Still closed due to winds. Took a pair of LUCI1/2 darks (dit x ndit = 3 x 20) for use in analyzing dither sequence. Set up dither sequence to dither NS rather than detxy.

05:00 Petr finished working on the 2019A version. Telescope is now running 2019A. ICE updates were deferred because of some problems running TCS complied with ICE 3.7 library. He notes in email that possible impacts are:
  • OSS UMACs controlling M2 and M3 support are running a new library for UMAC calls.
  • some minor updates to ECS
  • pmac code changed (for OSS PMACs)

06:44 Winds still gusting above 20m/s limit. Just had another one of 22 m/s at 06:53.

06:53 Shane starting to reconfigure for LN.

-- BarryRothberg - 25 May 2019
Topic revision: r6 - 27 May 2019, OlgaKuhn
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback